thermodynamic exercise:

When 1.0 g of fructose, C6H12O6 (s), a sugar commonly found in fruits is burned with oxygen in a calorimetric pump, the calorimeter temperature increases by 1.58 ° C. If the calorimetric capacity of the calorimetric pump (capacity 2 L) is 9.90 kJ / ° C. With the above information and some additional data determine the following:

A. The heat of molar combustion of fructose. What does the magnitude and sign indicate?
B. The enthalpy of molar combustion for fructose, if complete combustion is performed. What does the magnitude and sign indicate?
C. Does the enthalpy of combustion of fructose coincide with that reported by thermodynamic tables? If the difference is considerable, indicate the possible cause of error between the values.
Answer #1

a) The heat produced by combustion is calculated:

q = Ccal * ΔT = 9.90 kJ / ° C * 1.58 ° C = 15.64 kJ

Burned moles are calculated:

n = m / MM = 1 g / 180 g / mol = 5.6x10 ^ -3 mol

The heat of molar combustion is calculated:

molar = q / n = 15.64 kJ / 5.6x10 ^ -3 mol = 2800 kJ / mol

The heat is positive since it is absorbed by the calorimeter.

b) The molar enthalpy change is calculated:

ΔH = - q molar = - 2800 kJ / mol

The enthalpy change is negative because it represents the heat released by combustion.

c) The heat of combustion of fructose is reflected in the literature as - 2812 kJ / mol. It is observed that the calculated value is very similar to that of the literature.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.